OpenStudy (anonymous):

Easy question: a bird travels at 171 km/hr. Suppose you were to shoot the bird with a gun that fires at 366 m/s. If you fire at the instant the bird is 30m overhead how many meters ahead of you must you shoot. Ignore gravity in this example.

OpenStudy (anonymous):

(total distance)^2 = (amount ahead)^2 + (30m)^2 (total distance)/(366 m/s) = (amount ahead)/(171 km/hr) (1m/s)=(3.6km/hr)

OpenStudy (anonymous):

eliminating km/hr and total distance, we get amount ahead = (30m)/sqrt((366*3.6/171)^2 - 1) or 3.927 meters

OpenStudy (sriram):

i am solving this sum using SI units. 171kmph =47.5m/s assume distance ahead=x meters Time taken for bird to reach there=x/47.5 Time taken for bullet to reach there=distance travelled by bullet/366 distance travellrd by bullet=[x ^{2}+30^2]^1/2 both the times are equal.

OpenStudy (anonymous):

171km/hr=47.5m/s (47.5t)^2+(30)^2=(366t)^2 solving for t and multiplying by 47.5m/s u will get the distance.USED BYE THE METHOD OF RIGHT ANGLED TRIANGLE

OpenStudy (anonymous):

Answer = 30.25 m? i mean the hypotenuse of right angled triangle formed is 30.25?

OpenStudy (anonymous):

Hi Sriram i worked out the same was you did......You got 30.25m?